OpenStudy (anonymous):

3/5 (x+2)=x−4 WHAT DO I DO!?

OpenStudy (amistre64):

id assume you solve for x :)

OpenStudy (anonymous):

yes, but how!? hahah

OpenStudy (amistre64):

first get rid of the paranthesis by multiplying that 3/5 thru it; can you write that up?

OpenStudy (amistre64):

what does this become: 3/5 (x+2) when you multiply it thru?

OpenStudy (amistre64):

if fractions intimidate you we can start by clearing them, but I got no idea how to read your mind yet :)

OpenStudy (anonymous):

haha yes, I defitently have no idea how to do this..

OpenStudy (amistre64):

3/5 (x+2)=x−4 ; lets clear the fraction by multiplying everything we see by 5 and see if it makes it easier 3(x+2)=5x−20 ; how about now?

OpenStudy (anonymous):

alright, I understand that! now how do i finish it up? thank you so much, i'm taking my regents tomorrow to get a better grade and this is really helping me!

OpenStudy (amistre64):

well, we have 2 options, both are good; we can either divide it all by 3; or distribute the 3 thru the (...). Lets try the /3 option 3(x+2)=5x−20 /3 /3 /3 --------------- x+2 = 5x/3 - 20/3 ; now we get all our xs to one side by -x -x -x --------------- 2 = 5x/3 - x - 20/3 ; now move that 20/3 to the other side as well +20/3 +20/3 --------------------- 20/3 + 2 = 5x/3 -x ; factor out the x 20/3 + 2 = x(5/3 -1) ; and divide off the (5/3 -1) /(5/3-1) /5/3 -1 ------------------- 20/3 +2 -------- = x 5/3 -1

OpenStudy (amistre64):

there are prolly easier ways to do this, but you really give me nothing to go off of :)

OpenStudy (amistre64):

multiply it the left by 3/3 to get: 20+6 26 ----- = ----- = 13 = x 5-3 2

OpenStudy (amistre64):

algebra is a process; and it can be as hard or as easy as you want it to be ... the process is the same regardless, its how you go about it thats important tho

OpenStudy (anonymous):

Thank you so much! you really helped me. i'm so glad I found this website. Wish me luck tomorrow! I got a 75 in this regents the last time I took it but i'm hoping to get a 90 or better because I want to get into NYU! :)

OpenStudy (amistre64):

one way to solve this is pretty straight forward: 3/5 (x+2) = x−4 ; multiply the 3/5 thru the (..) 3x/5 + 6/5 = x-4 ; clear the fractions, *5 3x + 6 = 5x-20 ; gather the xs to one side 6 = 5x -3x -20 ; and the constants to the other 26 = 2x ; divide off the 4 26/2 = x ; and simplify 13 = x

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

thank you so much! :)